OpenStudy (anonymous):

solution of this question: SinA+SinB=1/4 & CosA+CosB=1/4 than Cos(A+B)=

OpenStudy (anonymous):

\[\sin A+\sin B=\frac{ 1 }{ 4 } \] \[2\sin \frac{ A+B }{ 2 }\cos \frac{ A-B }{ 2 }=\frac{ 1 }{ 4 } ...(1)\] \[\cos A+\cos B=\frac{ 1 }{ 4 }\] \[2\cos \frac{ A+B }{ 2 } \cos \frac{ A-B }{ 2 }=\frac{ 1 }{ 4 } ...(2)\] divide (1)by (2) \[\tan \frac{ A+B }{ 2 }=1=\tan \frac{ \pi }{ 4 }\] \[A+B=\frac{ \pi }{ 2 }\] \[\cos \left( A+B \right)=0\]